OpenStudy (anonymous):

simplify this expression ? (2^-1)^-3

OpenStudy (bibby):

\(\huge a^{-x}=\frac{1}{a^x}\)

OpenStudy (anonymous):

i think this one is more usefull: \( \Large (n^q)^p = n^{q*p}\)

OpenStudy (bibby):

oh I misread that as 2x-1 somehow

OpenStudy (anonymous):

np

OpenStudy (zale101):

2^-1=1/2^1 (1/2)^-3 1/1/(2^3) 1/1/8=8

OpenStudy (zale101):